# A Hybrid Data-Driven Multi-Stage Deep Learning Framework for Enhanced Nuclear Reactor Power Prediction

### v3

@@ Abstract
The accurate and efficient modeling of nuclear reactor transients is crucial for ensuring safe and optimal reactor operation. Traditional physics-based models, while valuable, can be computationally intensive and may not fully capture the complexities of real-world reactor behavior. This paper introduces a novel multi-stage deep learning framework that addresses these limitations, offering a faster and more robust solution for predicting the final steady-state power of reactor transients. By leveraging a combination of feed-forward neural networks with both classification and regression stages, and training on a unique dataset that integrates real-world measurements of reactor power and controls state from the Missouri University of Science and Technology Reactor (MSTR) with noise-enhanced simulated data, our approach achieves remarkable accuracy (96% classification, 2.3% MAPE). The incorporation of simulated data with noise significantly improves the model’s generalization capabilities, mitigating the risk of overfitting. This innovative solution not only enables rapid and precise prediction of reactor behavior but also has the potential to revolutionize nuclear reactor operations, facilitating enhanced safety protocols, optimized performance, and streamlined decision-making processes.
@@ 1 Introduction
Modeling is often used to monitor physical changes in a system or to determine the limitations of a design before it is implemented. Modeling software such as SCALE or TRACE has been specifically developed to model nuclear systems reliably without having to produce the system physically and allow for data collection at locations where sensors could not be feasibly placed in a real system. While these types of computational models have a high degree of accuracy, they are difficult to develop, tend to be system-specific, and are computationally intensive Samal et al. (2020). More specifically, without using significant computational resources, these models have difficulty making predictions with the desired degree of accuracy in the amounts of time required to make some operational decisions. These types of models, while they have their drawbacks, are still indispensable to nuclear engineers and designers due to their physics basing and high-fidelity nature.
@@
Due to recent developments in computing techniques and resources, new types of models have been utilized across all engineering disciplines. Deep Neural Networks (DNNs) are machine learning models that have received significant attention in recent years due to their versatility and well-documented development Massaoudi et al. (2021). DNNs have a variety of different internal architectures which have independent strengths and weaknesses and are applied in different instances depending on the problem being addressed Demuth et al. (2014). These types of models were designed using brains as a model, in which each neuron is connected to another with a certain weight which can then be adjusted to determine the strength of the connection. The result is a data-driven approach in which the DNN fits an input to an output by feeding the information through the neurons, which are collected in a series of layers Demuth et al. (2014). DNNs are beneficial compared to other types of models due to their relatively low consumption of computing resources after training and their ability to predict complicated phenomena without having to model each specific physical aspect in the system Lu et al. (2021).
@@
While DNNs are practical, they often avoid actually modeling the physics of the desired problem. These types of models typically function as direct input-to-output mapping versus true physics-based or probabilistic analysis Leva et al. (2017). That is, DNNs are more analogous to empirical models or equations with the input variables functioning as an initial condition. Physics-based models derive equations from known phenomena, whereas DNN models approximate an output value based on what it has learned from examining previous data points usually by fitting to a distribution Bin et al. (2013). As a practical model, DNNs are serviceable for accurate predictions. However, they are not a substitute for physical models if the desire is to understand how the system itself behaves. Furthermore, in many cases DNNs are trained using data collected or generated from physics-based models Kim et al. (2014). Furthermore, previous knowledge can be leveraged to produce models through processes such as transfer learning or multi-fidelity modeling which can infuse known physics information into machine learning models Pan (2023); Chen et al. (2021); Kabir et al. (2024).
@@
Similar to some other types of models in nuclear engineering, the DNNs are numerical models usually developed for a singular purpose. For a given DNN model, an input and output are defined before the model is developed. This generally results in a model that performs one specific task Sola et al. (1997). In the nuclear industry, the primary utility of DNN models is for safety-based predictions or thermal-hydraulic system analysis Bin et al. (2013). These areas of study are expected due to the design basis of nuclear reactors focusing on Loss of Coolant Accidents (LOCA) Kim et al. (2014); Do Koo et al. (2018). Additionally, the thermal-hydraulic system in a nuclear power plant is vital to both the safety and performance of the reactor. DNNs serve as excellent tools for determining the current state of flows or components in a thermal-hydraulic system which might be difficult to examine otherwise Radaideh et al. (2020).
@@
Predictive models are useful in nuclear systems for both operation and development, as demonstrated in our previous works Kobayashi et al. (2024a, b, c). DNNs serve as a valid option for each of these purposes for separate reasons. As an operational model, DNNs use minimal computing resources once trained and would allow for quick predictions within the desired degree of accuracy El-Sefy et al. (2021); Griesemer et al. (2023). This would allow operators to obtain additional information during an incident or to predict the system state ahead of time, each of which would allow for quicker reactions and a more detailed knowledge of the system at the desired time. Predictive model applications often have safety concepts implicitly built in as well. While not specifically aimed to address a safety problem, additional information, efficient automation procedure, and improved operator interfacing make systems and procedures safer Corrado (2021).
DNNs could also be useful for developing a full model by compiling multiple physical phenomena into a single input-output map. With the multiple different processes in a nuclear system affecting the system as a whole, the interactions between these processes can be difficult to model via physics-based equations Kim et al. (1993); Pickering et al. (2022). DNNs could allow for each of these phenomena to contribute to the output and model the effects of their interactions numerically without having to develop a complicated physical model Myung-Sub et al. (1991); Zhang et al. (2019); Pickering et al. (2022). In some disciplines, significantly robust models can be used to create a digital twin of a system. A DNN could be feasibly used to remedy this by simplifying a complicated multi-physics approach when used in conjunction with other models Shouman et al. (2022); Chen et al. (2021).
@@
The objectives of the study “Multi-Stage Deep Learning Model Development for State-Based Reactor Power Prediction" are focused on enhancing the prediction of nuclear reactor behaviors through the development and implementation of a novel deep learning approach. Specifically, the paper aims to:
@@
  • Combine Real-World and Simulated Data: Utilize a unique dataset that combines real-world operational data from the Missouri University of Science and Technology Reactor (MSTR) with simulated, noise-enhanced data. This approach aims to tackle common issues such as overfitting and to enhance the model’s robustness and ability to accurately reflect real-world reactor behaviors.
@@
  • Predict Reactor Power Output with High Accuracy: Ultimately, the objective is to efficiently predict reactor behavior, specifically the ending steady-state power output following reactor transients, with high levels of accuracy. The study seeks to demonstrate that the proposed multi-stage DL framework can achieve superior performance in this regard, offering a promising alternative for nuclear engineering applications.
@@
These objectives collectively target the advancement of nuclear reactor modeling and operational safety through the application of cutting-edge deep learning techniques, addressing existing limitations and opening new avenues for research and application in the field.
In detail, the multi-stage modeling approach is developed to improve on the standard feed-forward design and leverage existing advancements in machine learning. This style of model is developed by producing a classification result from a front-end model, which is then provided along with the input data to produce a regression prediction in the back-end model. This avoids complications associated with multi-task learning while still leveraging the benefits. This structure is chosen due to some of the prevalent modern strategies present in the machine learning research including:
  • Utilize transfer learning concepts by allowing a front-end model to generate a probabilistic prediction before a true (regression) prediction is made. While not directly utilizing transfer learning, the introduction of a predicted probabilistic input vector further informs a regression stage utilizing observed data.
  • Forced generalization using a probabilistic classification input vector, allowing the regression model to retrieve information about the local scope of the desired prediction. Due to the probabilistic nature of the classification vector, the regression layers are able to examine the scope in which the current prediction is being performed.
  • Pseudo-autoencoded input structure of models allows for initial information extraction, which can help achieve greater value from model inputs. This allow sthe model to process similar and relevant data before being used to directly make a prediction in later layers.
These strategies can be combined to result in a computationally cheap model which can be leveraged for decision-making or short-term modeling. The goal of this model is to provide information to operators and engineers in a timely manner to improve operation efficiency and allow for additional, readable information to be provided to personnel.
@@
  • Multi-Stage Deep Learning Framework: A novel framework that leverages a combination of feed-forward neural networks for both classification and regression, specifically tailored for predicting reactor power states. This is a significant departure from single-stage models, allowing for more nuanced analysis and predictions.
  • Hybrid Data Approach: The method of combining real-world measurements with simulated, noise-enhanced data for training the model is innovative. It not only addresses the challenge of dataset size but also improves the model’s ability to predict real-world reactor behavior accurately.
@@ 2 Procedure and Multi-Stage DNN Development
Developing a deep learning model consists of developing a pipeline for the data, training, and model development. The following section outlines the development of the pipeline in the standard order, specifically data collection, data processing, and model development. The training and validation results are excluded from this section and instead included in the Results and Discussion.
@@ 2.1 Data Collection
MSTR is the research nuclear reactor on the campus of Missouri S&T. This reactor, with a maximum rated power of 200kW, is used frequently for training reactor operators and other personnel compared to other research reactors, such as those focusing on irradiation experiments. This presents a unique data space compared to the other reactors, resulting in frequent power changes. This data space can then be utilized to extract additional information regarding power changes in small reactor systems in general. The bulk of power change data available from MSTR includes power changes from operator training and reactor operations courses. This is collected as state-based data, with each state including control rod heights and reactor power.
Data was collected from MSTR via daily transient logs as shown in Figure 1. A single power transient could be defined by generating a start point and end point based on stable reactor power. These two points in time could then be used to define a state-based approach to the reactor transient, where some information from the initial and final state could be used in conjunction to make predictions.
@@
The same information was collected from each stable power, with two stable powers forming a single observation for the state-based prediction model. The information from any given stable power could be coupled with the previous or subsequent information from a stable power to create a single data point. The information collected from each of these points included the stable power, control rod heights, date, and time.
Data could be collected in this format and exported to an Excel file. Each observation included the date, start time, end time, initial power, final power, and initial and final control rod heights. Date and time data were retained for indexing, while the power and control rod information could be used to define the physical state of the reactor system. Transient data ranging from January 7, 2013, to October 10, 2014, was collected and processed by hand in this manner.
Some observations were excluded from the collected data. For example, reactor shutdowns were not included due to the excess negative reactivity in the core, which could result in model memorization of shutdowns or negatively impact performance by preferential training towards shutdowns. Transients longer than one hour (from start to finish) were excluded to focus the model on short-term state-based modeling. After data auditing and exclusion, 542 total observations were collected.
Unfortunately, 542 observations would not be enough to train a model adequately. To address the challenge associated with the inadequate training data, a physical sampling method was implemented. The power defect equation, a generalization of the prompt jump equation, was used to generate these additional samples Lamarsh et al. (2001). This equation is generalized as equation 1. The power defect was applied to the final state in observation to jump from one stable power to another stable power effectively. The power defect method was performed under the condition that the total reactivity change in this state could not exceed 0.5\$ and no control rod exceeded the 24 inches of maximum withdrawal. This sampling method could be applied $n$ times and form the bulk of the training data while retaining the actual observations for the testing set. The relationship between the reactor power and reactivity can be expressed by Eq. 1:
@@
A method was defined using a Python code to randomly perturb control rod heights from both the initial and final states of observation. The functions used to execute this practice are included as 5.5 and LABEL:sec:over_sample_func. Both the initial and final states were perturbed with this method in an effort to reduce the memorization capabilities of the model. The reactivity was determined by using the integral rod worth information for the relevant core configuration based on the date the transient occurred. Over the two years of training samples collected, four different reactor configurations were utilized. The date of the observation could be cross-referenced to the operational dates of the core configurations in order to determine which set of control rod-worth values should be used for the selected sample. For the conversion of the given rod worth values, a delayed neutron fraction ($\beta$) of 0.006 was assumed.
An additional dataset generation method using a joint statistical distribution was performed to increase dataset breadth and smoothness. A random normal value was generated, which then served as the upper and lower limit to a random normal distribution. This random normal distribution was used to perturb control rod heights and then calculate reactivity insertion. This methodology increases diversity in individual control rod heights and reduces direct clustering around semi-random points in the data. Additional dataset generation methods may be provided upon request to authors.
Table 1 represents the relationship between the core configuration and the control rod worths. Using the control rod worths, additional data points could be generated using some random noise. The original dataset and generated dataset can be seen in Figure 2.

### v4

# Digital Twin-Centered Hybrid Data-Driven Multi-Stage Deep Learning Framework for Enhanced Nuclear Reactor Power Prediction

@@ Abstract
The accurate and efficient modeling of nuclear reactor transients is crucial for ensuring safe and optimal reactor operation. Traditional physics-based models, while valuable, can be computationally intensive and may not fully capture the complexities of real-world reactor behavior. This paper introduces a novel hybrid digital twin-focused multi-stage deep learning framework that addresses these limitations, offering a faster and more robust solution for predicting the final steady-state power of reactor transients. By leveraging a combination of feed-forward neural networks with both classification and regression stages, and training on a unique dataset that integrates real-world measurements of reactor power and controls state from the Missouri University of Science and Technology Reactor (MSTR) with noise-enhanced simulated data, our approach achieves remarkable accuracy (96% classification, 2.3% MAPE). The incorporation of simulated data with noise significantly improves the model’s generalization capabilities, mitigating the risk of overfitting. Designed as a digital twin supporting system, this framework integrates real-time, synchronized predictions of reactor state transitions, enabling dynamic operational monitoring and optimization. This innovative solution not only enables rapid and precise prediction of reactor behavior but also has the potential to revolutionize nuclear reactor operations, facilitating enhanced safety protocols, optimized performance, and streamlined decision-making processes. By aligning data-driven insights with the principles of digital twins, this work lays the groundwork for adaptable and scalable solutions in nuclear system management.
@@ 1 Introduction
Modeling is often used to monitor physical changes in a system or to determine the limitations of a design before it is implemented. Modeling software such as SCALE or TRACE has been specifically developed to model nuclear systems reliably without having to produce the system physically and allow for data collection at locations where sensors could not be feasibly placed in a real system. While these types of computational models have a high degree of accuracy, they are difficult to develop, tend to be system-specific, and are computationally intensive Samal et al. (2020). More specifically, without using significant computational resources, these models have difficulty making predictions with the desired degree of accuracy in the amounts of time required to make some operational decisions. These types of models, while they have their drawbacks, are still indispensable to nuclear engineers and designers due to their physics-basing and high-fidelity nature.
@@
Due to recent developments in computing techniques and resources, new types of models have been utilized across all engineering disciplines. Deep Neural Networks (DNNs) are machine learning models that have received significant attention in recent years due to their versatility and well-documented development Massaoudi et al. (2021). DNNs have a variety of different internal architectures, which have independent strengths and weaknesses and are applied in different instances depending on the problem being addressed Demuth et al. (2014). These types of models were designed using brains as a model, in which each neuron is connected to another with a certain weight which can then be adjusted to determine the strength of the connection. The result is a data-driven approach in which the DNN fits an input to an output by feeding the information through the neurons, which are collected in a series of layers Demuth et al. (2014). DNNs are beneficial compared to other types of models due to their relatively low consumption of computing resources after training and their ability to predict complicated phenomena without having to model each specific physical aspect in the system Lu et al. (2021).
@@
While DNNs are practical, they often avoid modeling the physics of the desired problem. These models typically function as direct input-to-output mapping versus true physics-based or probabilistic analysis Leva et al. (2017). That is, DNNs are more analogous to empirical models or equations with the input variables functioning as an initial condition. Physics-based models derive equations from known phenomena, whereas DNN models approximate an output value based on what it has learned from examining previous data points, usually by fitting to a distribution Bin et al. (2013). To function as a true digital twin Yadav et al. (2023); Liu et al. (2024); Kobayashi et al. (2024a, b, c), a model must integrate the strengths of data-driven methods with physics-informed insights, enabling it to not only predict system behavior but also adapt dynamically to operational changes in real-time. Furthermore, in many cases, DNNs are trained using data collected or generated from physics-based models Kim et al. (2014). Furthermore, previous knowledge can be leveraged to produce models through processes such as transfer learning or multi-fidelity modeling, which can infuse known physics information into machine learning models Pan (2023); Chen et al. (2021); Kabir et al. (2024).
@@
Similar to other types of nuclear engineering models, the DNNs are numerical models usually developed for a singular purpose. For a given DNN model, an input and output are defined before the model is developed. This generally results in a model that performs one specific task Sola et al. (1997). In the nuclear industry, the primary utility of DNN models is for safety-based predictions or thermal-hydraulic system analysis Bin et al. (2013). These areas of study are expected due to the design basis of nuclear reactors focusing on Loss of Coolant Accidents (LOCA) Kim et al. (2014); Do Koo et al. (2018). Additionally, the thermal-hydraulic system in a nuclear power plant is vital to both the safety and performance of the reactor. DNNs serve as excellent tools for determining the current state of flows or components in a thermal-hydraulic system, which might be difficult to examine otherwise Radaideh et al. (2020). When adapted to digital twin frameworks and systems Yadav et al. (2023), these models can provide real-time insights into reactor state, enabling predictive diagnostics, anomaly detection, and operational optimization.
@@
Predictive models are useful in nuclear systems for both operation and development, as demonstrated in our previous works Kobayashi et al. (2024a, b, c). DNNs serve as a valid option for each of these purposes for separate reasons. As the foundation of a digital twin Yadav et al. (2023); Liu et al. (2024), DNNs can act as a real-time decision support system, offering operators immediate insights into reactor states and predictive analysis during incidents. By leveraging their minimal computational resource requirements post-training, these models can make fast predictions within the desired degree of accuracy El-Sefy et al. (2021); Griesemer et al. (2023). This would allow operators to obtain additional information during an incident or to predict the system state ahead of time, each of which would allow for quicker reactions and a more detailed knowledge of the system at the desired time. Predictive model applications often have safety concepts implicitly built in as well. While not specifically aimed to address a safety problem, additional information, efficient automation procedure, and improved operator interfacing make systems and procedures safer Corrado (2021). Combining these features into a synchronized framework, a digital twin offers a transformational leap in monitoring, control, and decision-making of nuclear reactor systems. Predictive models are useful in nuclear systems for both operation and development, as demonstrated in previous works Kobayashi et al. (2024a, b, c). DNNs serve as a valid option for each of these purposes for separate reasons. As the foundation of a digital twin Yadav et al. (2023); Liu et al. (2024), DNNs can act as a real-time decision support system, offering operators immediate insights into reactor states and predictive analysis during incidents. By leveraging their minimal computational resource requirements post-training, these models can make fast predictions within the desired degree of accuracy El-Sefy et al. (2021); Griesemer et al. (2023). This would allow operators to obtain additional information during an incident or to predict the system state ahead of time, each of which would allow for quicker reactions and a more detailed knowledge of the system at the desired time. Predictive model applications often have safety concepts implicitly built in as well. While not specifically aimed to address a safety problem, additional information, efficient automation procedure, and improved operator interfacing make systems and procedures safer Corrado (2021). By combining these features into a synchronized framework, a digital twin offers a transformational leap in nuclear reactor monitoring and control. Predictive models are useful in nuclear systems for both operation and development, as demonstrated in our previous works Kobayashi et al. (2024a, b, c). DNNs serve as a valid option for each of these purposes for separate reasons. As the foundation of a digital twin Yadav et al. (2023); Liu et al. (2024), DNNs can act as a real-time decision support system, offering operators immediate insights into reactor states and predictive analysis during incidents. By leveraging their minimal computational resource requirements post-training, these models can make fast predictions within the desired degree of accuracy El-Sefy et al. (2021); Griesemer et al. (2023). This would allow operators to obtain additional information during an incident or to predict the system state ahead of time, each of which would allow for quicker reactions and a more detailed knowledge of the system at the desired time. Predictive model applications often have safety concepts implicitly built in as well. While not specifically aimed to address a safety problem, additional information, efficient automation procedure, and improved operator interfacing make systems and procedures safer Corrado (2021). By combining these features into a synchronized framework, a digital twin offers a transformational leap in nuclear reactor monitoring and control. Predictive models are useful in nuclear systems for both operation and development, as demonstrated in our previous works Kobayashi et al. (2024a, b, c). DNNs serve as a valid option for each of these purposes for separate reasons. As the foundation of a digital twin Yadav et al. (2023); Liu et al. (2024), DNNs can act as a real-time decision support system, offering operators immediate insights into reactor states and predictive analysis during incidents. By leveraging their minimal computational resource requirements post-training, these models can make fast predictions within the desired degree of accuracy El-Sefy et al. (2021); Griesemer et al. (2023). This would allow operators to obtain additional information during an incident or to predict the system state ahead of time, each of which would allow for quicker reactions and a more detailed knowledge of the system at the desired time. Predictive model applications often have safety concepts implicitly built in as well. While not specifically aimed to address a safety problem, additional information, efficient automation procedure, and improved operator interfacing make systems and procedures safer Corrado (2021). By combining these features into a synchronized framework, a digital twin offers a transformational leap in nuclear reactor monitoring and control. DNNs could also be useful for developing a full model by compiling multiple physical phenomena into a single input-output map. With the multiple different processes in a nuclear system affecting the system as a whole, the interactions between these processes can be difficult to model via physics-based equations Kim et al. (1993); Pickering et al. (2022). Such an approach aligns with the principles of digital twins Yadav et al. (2023) by enabling real-time, system-wide predictions while remaining adaptable to changing conditions. DNNs could allow for each of these phenomena to contribute to the output and model the effects of their interactions numerically without having to develop a complicated physical model Myung-Sub et al. (1991); Zhang et al. (2019); Pickering et al. (2022). In some disciplines, significantly robust models can be used to create a digital twin of a system Yadav et al. (2023). A DNN could be feasibly used to remedy this by simplifying a complicated multi-physics approach when used in conjunction with other models Shouman et al. (2022); Chen et al. (2021).
@@
The objectives of the study are focused on enhancing the prediction of nuclear reactor behaviors through the development and implementation of a novel deep learning approach. Specifically, the paper aims to:
@@
  • Combine Real-World and Simulated Data: Utilize a unique dataset that combines real-world operational data from the Missouri University of Science and Technology Reactor (MSTR) with simulated, noise-enhanced data. This approach aims to tackle common issues such as overfitting and to enhance the model’s robustness and ability to reflect real-world reactor behaviors accurately.
@@
  • Predict Reactor Power Output with High Accuracy: Ultimately, the objective is to predict reactor behavior efficiently, specifically the ending steady-state power output following reactor transients, with high levels of accuracy. The study seeks to demonstrate that the proposed multi-stage DL framework can achieve superior performance in this regard, offering a promising alternative for nuclear engineering applications.
@@
These objectives collectively target the advancement of nuclear reactor modeling and operational safety through the application of cutting-edge deep learning techniques. By addressing existing limitations and incorporating real-time predictive and prescriptive capabilities, this work lays the foundation for the development of a machine learning framework Yadav et al. (2023) for digital twin-based reactor operations, opening new avenues for research and application in the field.
In detail, the multi-stage modeling approach is developed to improve on the standard feed-forward design and leverage existing advancements in machine learning, with a clear alignment to digital twin principles. This style of model is developed by producing a classification result from a front-end model, which is then provided along with the input data to produce a regression prediction in the back-end model. This avoids complications associated with multi-task learning while still leveraging the benefits. Such a structure is critical for ensuring the modularity, real-time inference, and adaptability expected from a digital twin. This structure is chosen due to some of the prevalent modern strategies present in machine learning research, including:
  • Utilize transfer learning concepts by allowing a front-end model to generate a probabilistic prediction before making a true (regression) prediction. While not directly utilizing transfer learning, the introduction of a predicted probabilistic input vector further informs a regression stage utilizing observed data.
  • The probabilistic classification input vector enables the regression model to adapt to local conditions, ensuring robustness and accuracy across varying reactor states. **This adaptability supports the digital twin’s requirement to operate reliably under diverse scenarios.
  • By extracting and processing relevant information early, the Pseudo-autoencoded input structure of models ensures computational efficiency and meaningful feature representation, a necessity for digital twins to perform real-time inference and diagnostics effectively.
These strategies combine to create a computationally efficient model that is deployable in a digital twin prototype and provides timely, actionable insights to operators and engineers. The model enhances operational efficiency by enabling informed decision-making and presenting additional, interpretable information to personnel in critical moments.
@@
  • Multi-Stage Deep Learning Framework: A novel framework combining feed-forward neural networks for classification and regression, specifically tailored for predicting reactor power states. This modular, layered design reflects the architecture of digital twins, facilitating synchronization with reactor operations and enabling predictive insights.
  • Hybrid Data Approach: Integrating real-world reactor measurements with simulated, noise-enhanced data for training ensures robustness and reliability. This hybrid strategy supports the production of a high-fidelity digital twin that can generalize across real-world reactor conditions and respond dynamically to changes.
@@ 2 Procedure and Multi-Stage DNN Development
Developing a deep learning model involves developing a pipeline for the data, training, and model development. The following section outlines the development of the pipeline in the standard order, specifically data collection, data processing, and model development. The training and validation results are excluded from this section and instead included in the Results and Discussion.
@@ 2.1 Data Collection
MSTR is the research nuclear reactor on the campus of Missouri S&T. This reactor, with a maximum rated power of 200kW, is used frequently for training reactor operators and other personnel compared to other research reactors, such as those focusing on irradiation experiments. This operational diversity provides a dynamic dataset well-suited for developing a digital twin capable of capturing frequent power changes and transient dynamics. This data space can then be utilized to extract additional information regarding power changes in small reactor systems in general. The bulk of power change data available from MSTR includes power changes from operator training and reactor operations courses. The state-based nature of this data aligns with digital twin requirements for accurate system representation and real-time synchronization.
Data was collected from MSTR via daily transient logs as shown in Figure 1. The transient logs provide a foundation for state-based digital twin modeling, where stable reactor power states can define each transient. A single power transient could be defined by generating start and end points based on stable reactor power. These two points in time could then be used to define a state-based approach to the reactor transient, where some information from the initial and final state could be used to make predictions. This state-based approach mirrors the system’s physical behavior and ensures a digital twin has the capability to remain synchronized with real-world operations.
@@
The same information was collected from each stable power, with two stable powers forming a single observation for the state-based prediction model. These stable power observations serve as discrete snapshots of the system’s operational state, forming the building blocks for the digital twin. The information from any given stable power could be coupled with the previous or subsequent information from a stable power to create a single data point. The information collected from each of these points included the stable power, control rod heights, date, and time. The digital twin can dynamically predict and simulate the system’s future states by retaining this information.
Data could be collected in this format and exported to an Excel file. Each observation included the date, start time, end time, initial power, final power, and initial and final control rod heights. Date and time data were retained for indexing, while the power and control rod information could be used to define the physical state of the reactor system. This comprehensive data format supports a digital twin’s requirement for high-fidelity synchronization with real-world operations. Transient data ranging from January 7, 2013, to October 10, 2014, was collected and processed by hand in this manner.
Some observations were excluded from the collected data. For example, reactor shutdowns were not included due to the excess negative reactivity in the core, which could result in model memorization of shutdowns or negatively impact performance by preferential training towards shutdowns. Transients longer than one hour (from start to finish) were excluded to focus the model on short-term state-based modeling. This focus on short-term transients aligns with the operational needs of a digital twin, ensuring accurate and timely predictions for operational decision-making. After data auditing and exclusion, 542 total observations were collected.
Unfortunately, 542 observations would not be enough to train a model adequately. A physical sampling method was implemented to address the challenge associated with the inadequate training data. The power defect equation, a generalization of the prompt jump equation, was used to generate these additional samples Lamarsh et al. (2001). This equation is generalized as equation 1. This augmentation process expands the model’s training dataset, enabling it to generalize across diverse operational scenarios while preserving fidelity to physical principles and enabling expanded deployment in digital twin-type systems. The power defect was applied to the final state in observation to jump from one stable power to another stable power effectively. The power defect method was performed under the condition that the total reactivity change in this state could not exceed 0.5\$ and no control rod exceeded the 24 inches of maximum withdrawal. This sampling method could be applied $n$ times and form the bulk of the training data while retaining the actual observations for the testing set. The relationship between the reactor power and reactivity can be expressed by Eq. 1:
@@
A method was defined using a Python code to randomly perturb control rod heights from both the initial and final states of observation. These perturbations add diversity to the dataset, enhancing the model’s robustness and adaptability to dynamic system changes. The functions used to execute this practice are included as 5.5 and 5.6. Both the initial and final states were perturbed with this method in an effort to reduce the memorization capabilities of the model. The reactivity was determined by using the integral rod worth information for the relevant core configuration based on the date the transient occurred. Over the two years of training samples collected, four different reactor configurations were utilized. The date of the observation could be cross-referenced to the operational dates of the core configurations in order to determine which set of control rod-worth values should be used for the selected sample. For the conversion of the given rod worth values, a delayed neutron fraction ($\beta$) of 0.006 was assumed.
An additional dataset generation method using a joint statistical distribution was performed to increase dataset breadth and smoothness. This method enhances the model’s ability to simulate diverse operational scenarios by generating data with statistical variability and smooth transitions between states, necessary for sufficiently robust operation in digital twins. A random normal value was generated, which then served as the upper and lower limit to a random normal distribution. This random normal distribution was used to perturb control rod heights and then calculate reactivity insertion. This methodology increases diversity in individual control rod heights and reduces direct clustering around semi-random points in the data. Additional dataset generation methods may be provided upon request to authors.
Table 1 represents the relationship between the core configuration and the control rod worths. Using the control rod worths, additional data points could be generated using some random noise. These enriched datasets ensure that the model reflects real-world variability while maintaining high fidelity to reactor physics. The original dataset and generated dataset can be seen in Figure 2.